OpenStudy (anonymous):

What is the value of 4C2?

OpenStudy (unklerhaukus):

six

OpenStudy (anonymous):

\[nCr = \frac{n!}{(n - r)! r!}\]

OpenStudy (anonymous):

4C2 \[=\frac{4!}{(4-2)!2!}\]\[=\frac{ 4!}{2!2!}\]\[ = 6\]

OpenStudy (anonymous):